OpenStudy (anonymous):

derivative of (x+1)^3(6-x)^2

OpenStudy (anonymous):

4566=jojo

OpenStudy (anonymous):

(uv)'=u'v+uv' in your case u=(x+1)^3, v=(6-x)^2

OpenStudy (anonymous):

First use the product rule, the two functions being (x+1)^3 and (6-x)^2 The derivative of these two functions must consequently be found by the chain rule. So if\[u=(x+1)^3 and, v=(6-x)^2\] Then by the chain rule, \[u'=3(x+1)^2, v'=-2(6-x)\] Given u' and v' you should be able to put it together via the product rule. as mentioned by myko.

OpenStudy (anonymous):

okay, thank you guys!